P5-43 Deposits to accumulate future sums For each case shown in the following tabie. determine the amount of the equal, end-of-year deposits necessary to the given sum at the end of the specified period, assuming the stated annual interest rate. accumulate Sum to bc Accumulation Interest Case accumulated pcriod (ycars) rate 12% $ 5,000 100,000 30,000 15,000 20 10 12

Please show how you got to this answer. Thank you!
Answer #1

Future value of annuity=Annuity[(1+rate)^time period-1]/rate

A:

5000=Annuity[(1.12)^3-1]/0.12

5000=Annuity*3.3744

Annuity=5000/3.3744

=$1481.74(Approx).

B:

100,000=Annuity[(1.07)^20-1]/0.07

100,000=Annuity*40.99549232

Annuity=100,000/40.99549232

=$2439.29(Approx).

C:

30,000=Annuity[(1.1)^8-1]/0.1

30,000=Annuity*11.4358881

Annuity=30,000/11.4358881

=$2623.32(Approx).

D:

15000=Annuity[(1.08)^12-1]/0.08

15000=Annuity*18.97712646

Annuity=15000/18.97712646

=$790.43(Approx).